The following functions pertain to the loanable funds market of a hypothetical country:
DL= 41 - 7r
SL = 5 + 2r
DL = SL
Where DL is demand for loanable funds, SL is supply of loanable funds and r is the real interest rate.
a) Find the equilibrium values of r, DL and SL.
b) If there is GHȼ81 million-bond financed increase in government expenditure, find:
i) The new equilibrium values of r, DL and SL.
1
Expert's answer
2020-08-24T14:58:37-0400

Solution:

a.) Equilibrium: DL = SL

"DL = SL"

"41-7r = 5+2r"

"41-5=2r+7r"

"36 = 9r"

"r=\frac{36}{9}"


"r = 4"


b.) An increase in government expenditure by GHȼ81 million-bond will lead to an increase in demand for loanable funds due to high interest rate incurred and an increase in supply of loanable funds.

New demand equation will be:

"41-81-7r = -40-7r"


New supply function will be:

"5+81+2r =86+2r"


New equilibrium will be:

DL = SL

"-40-7r=86+2r"

"-40-86=2r+7r"

"-126=9r"

"r=\frac{-126}{9}"


"r=-14"